suppose that a judge's decision follows a binomial distribution and that his verdict is correct 90% of the time. in his next 10 decisions, what is the probability that he makes exactly 8 correct decisions?
Answer by sudhanshu_kmr(1152)


probability of correct decision = 90/100= 9/10
probability of incorrect decision = 1 - 9/10 = 1/10

probability of exactly 8 correct decision out of 10 = 10C8*(9/10)^8 *(1/10)^2

= 45*(9/10)^8 *(1/10)^2